OpenStudy (anonymous):

Solve the following system using the substitution method. 3x – y = –2 6x + 5y = 24

OpenStudy (anonymous):

y=3x+2 Now got to the equation 6x+5y=24 and subs. instead of y .. 6x+5(3x+2)=24 and solve

OpenStudy (anonymous):

y = 3x + 2 6x + 15x + 10 = 24 21x = 14 x = 2/3 and now find y..

OpenStudy (anonymous):